### v2

# Relic gravitational waves from primordial gravitational collapses

Abstract
A large primordial density perturbation of the Hubble scale will gravitationally collapse, generating an outgoing sound shell whether or not a primordial black hole (PBH) is formed. In this Letter, we report a new source of the stochastic gravitational wave background induced by the collision of sound shells in the early Universe. The peak frequency and amplitude in the GW spectrum depend on the Hubble horizon and the abundance of sound shells. Abundant density perturbations would lead to GW backgrounds potentially detectable for future pulsar timing arrays and ground-based/space-borne detectors. For those perturbations that collapse into PBHs, future null detection of the corresponding high-frequency GW background could put new observational constraints on those PBHs that have already evaporated.
Introduction. Recently, various Pulsar Timing Array (PTA) collaborations (NANOGrav Agazie et al. (2023), EPTA Antoniadis et al. (2023), PPTA Reardon et al. (2023), and CPTA Xu et al. (2023)) have announced possible detections of stochastic gravitational-wave backgrounds (SGWBs). With more upcoming GW detectors (such as space-borne detectors like LISA Amaro-Seoane et al. (2017), Taiji Hu and Wu (2017); Ruan et al. (2020), and TianQin Luo et al. (2016), as well as the more advanced DECIGO Kawamura et al. (2006) and BBO Corbin and Cornish (2006), along with ground-based detectors like Einstein Telescope (ET) Maggiore et al. (2020) and Cosmic Explorer (CE) Reitze et al. (2019)), we might embrace many SGWBs Caprini and Figueroa (2018) with rich new physics Cai et al. (2017); Bian et al. (2021). Therefore, it is both observationally and theoretically appealing to identify a new SGWB.
It has long been expected Hawke and Stewart (2002); Musco et al. (2005); Musco and Miller (2013); Bloomfield et al. (2015); García-Bellido et al. (2021); Carr et al. (2021) that an outgoing sound shell Deng and Vilenkin (2017); Deng et al. (2018); Deng (2020a, b) will form along with primordial black hole (PBH) formation from gravitational collapse of a sufficiently large density perturbation of horizon size. In this Letter, we will report a new source of SGWBs induced by the superposition of sound waves from gravitational collapses of large density perturbations of horizon size in the early Universe via full numerical relativity. The details of numerical simulations are presented separately in an accompanying long paper Ning et al. (2025) on the long-time evolution of such large density perturbations. This Letter will focus on the GW phenomenology even if these density perturbations are not large enough to form PBHs.
Sound waves from density perturbations. PBHs can form in the early Universe if the perturbation amplitude exceeds a critical threshold Carr and Hawking (1974); Hawking (1971). For those density perturbations of horizon size that are not sufficiently dense to result in PBH formation, bounces occur, and a shell of density contrast dubbed sound wave propagates outwards, as shown in previous numerical simulation results Hawke and Stewart (2002); Musco et al. (2005); Bloomfield et al. (2015); Musco and Miller (2013). In the accompanying long paper Ning et al. (2025), we find that the sound wave could also form for the case with PBH formation, and hence, we will follow the numerical treatment of the PBH formation.
Figure 1: Evolution of the energy density in the process of PBH formation. The upper panel describes the sub-critical case with $\delta_{m}=0.3$. In this case, only a sound wave is formed, as the perturbation is not dense enough. The middle panel describes the near-critical case with $\delta_{m}=0.5$. Due to the large gradient of fluid pressure, both underdense and overdense structures appear. The lower panel denotes the super-critical case with $\delta_{m}=0.55$. In this case, there is only an underdense structure.
Figure 2: The density contrasts $\delta$ (top row) and the comoving radial velocity $u$ (bottom row) of the fluid for the sub-critical case with $\delta_{m}=0.3$ (left column), near-critical case with $\delta_{m}=0.5$ (middle column), and super-critical case with $\delta_{m}=0.55$ (right column), where the PBHs are formed around $t=26t_{m}$ and $t=8.6t_{m}$ with the initial PBH comoving radius $r_{\mathrm{PBH}}=0.59r_{H}$ and $r_{\mathrm{PBH}}=1.97r_{H}$, respectively.
Assuming spherical symmetry, numerical simulation conventionally solves the Misner-Sharp equations Misner and Sharp (1964) for the motion of the relativistic fluid with a perfect fluid ansatz for the energy-momentum tensor $T^{\mu\nu}=(p+\rho)u^{\mu}u^{\nu}+pg^{\mu\nu}$, where $u^{\mu}$ is the fluid four-velocity and $p$ and $\rho$ are related by an equation-of-state parameter $\omega\equiv p/\rho$. The line element with spherical symmetry reads $\mathrm{d}s^{2}=-A(r,t)^{2}\mathrm{d}t^{2}+B(r,t)^{2}\mathrm{d}r^{2}+R(r,t)^{2}\mathrm{d}\Omega^{2}$ with $\mathrm{d}\Omega^{2}=\mathrm{d}\theta^{2}+\mathrm{sin}^{2}(\theta)\mathrm{d}\phi^{2}$. At superhorizon scales, the metric can be approximated at the leading order of gradient expansion as Shibata and Sasaki (1999)
$\mathrm{d}s^{2}=-\mathrm{d}t^{2}+a(t)^{2}\left[\frac{\mathrm{d}r^{2}}{1-K(r)r^{2}}+r^{2}\mathrm{d}\Omega^{2}\right],$ (1)
where the initial curvature $K(r)$ encodes the information of curvature perturbations Hidalgo and Polnarev (2009); Musco (2019) that further induce density perturbations. In this Letter, we simply follow the conventional treatment by adopting a Gaussian profile for the initial curvature perturbation,
$K(r)=\mathcal{A}e^{-(r/r_{m})^{2}},$ (2)
where $r_{m}$ is the length scale of the perturbation and $\mathcal{A}$ denotes its peak amplitude. The curvature perturbation can induce density contrast $\delta_{m}\equiv\delta(r_{m},t_{m})$ at the horizon crossing time $t_{m}$ via $\delta_{m}=\frac{3(1+\omega)}{e(5+3\omega)}\mathcal{A}r_{m}^{2}$ Musco (2019). With the above initial profile, the threshold for PBH formation is $\delta_{c}\simeq 0.4977$ in the radiation-dominated (RD) era with $\omega=1/3$ as obtained in our accompanying long paper Ning et al. (2025), which numerically evolves the density perturbation collapse for the first time in the long-time limit. We modify the code developed in Ref. Escrivà (2020) and perform all of our simulations in the RD era, but the generalization to other $\omega$ should be straightforward.
Some typical snapshots of our simulations with initial condition $r_{m}=10r_{H}$ ($r_{H}$ is the initial Hubble horizon) are shown in Fig. 1 from top down for the sub-critical ($\delta_{m}=0.3$), near-critical ($\delta_{m}=0.5$), and super-critical ($\delta_{m}=0.55$) collapse cases. The corresponding radial profiles are shown in Fig. 2 for the density contrast (top row) and comoving velocity normalized by the background fluid velocity (bottom row) in the sub-critical (left column), near-critical (middle column), and super-critical (right column) cases, respectively. Both quantities are expressed in terms of comoving radial coordinates. The shape of the sound shell depends on the amplitude of the perturbation $\delta_{m}$ at horizon reentry. A sub-critical or near-critical case will lead to sound waves with both overdense and underdense shells, for the fluid pressure gradient is large enough to push the fluid away. A super-critical case will only admit an underdense shell.
Figure 3: Left: The dimensionless GW power spectrum with an inset to the infrared end. Right: The present GW energy density spectrum. We have used Schmitz to plot the sensitivity curves. In both panels, the dot-dash line denotes the sub-critical case $\delta_{m}=0.3$ with collision time $t_{c}=50t_{m},100t_{m}$, while the solid line denotes the near-critical case $\delta_{m}=0.5$ with collision time $t_{c}=200t_{m},400t_{m}$, and the dashed line denotes the super-critical case $\delta_{m}=0.55$ with collision time $t_{c}=400t_{m},800t_{m}$.
GW spectra from sound waves. To estimate the GW spectrum, we closely follow the sound shell model Hindmarsh (2018); Hindmarsh and Hijazi (2019); Guo et al. (2021) (see also Refs. Cai et al. (2023); Roper Pol et al. (2024); Sharma et al. (2023) for recovering the correct infrared behavior Cai et al. (2020)) to adapt to our case as detailed in the Supplementary Material, where the only difference is an input velocity power spectrum that sources the energy of GWs. For the case of a first-order phase transition (FOPT) after initial acceleration stage Cai and Wang (2021), the fluid velocity profile is self-similar Espinosa et al. (2010) driven by a constantly expanding wall until wall collision; from then on, the sound shell propagates freely with its velocity profile at the time as the input in the sound shell model.
In our case, there is no bubble wall pushing outward the sound shell so that we can directly extract from our numerical simulations the velocity profile just a moment before the sound shell collision (to make the sound shell propagate as freely as possible) as the input in the sound shell model, and derive the fluid velocity power spectrum,
$\langle v_{\vec{q}}^{i}v_{\vec{k}}^{j*}\rangle=\sum_{m=1}^{N}\sum_{n=1}^{N}\langle\hat{q}^{i}\hat{k}^{j}A(q)A^{*}(k)e^{-i\vec{q}\cdot\vec{x}^{(m)}+i\vec{k}\cdot\vec{x}^{(n)}}e^{i(\omega_{q}-\omega_{k})t_{i}}\rangle,$ (3)
where $\omega_{q}\equiv c_{s}q=q/\sqrt{3}$, $N$ is the number of sound shells at a given spatial point with their center positions labeled by $\vec{x}^{(m)}$, $t_{i}$ denotes the time when the sound shell starts to propagate freely  [^1] This formalism of fluid velocity power spectrum implies that we have assumed all sound shells are formed roughly at the same time, which is the case for the monochromatic PBH formations. and $A(q)$ encodes the information of the initial profile of density contrast and fluid velocity  [^2] For a more accurate definition, one refers to Ref. Roper Pol et al. (2024). Nevertheless, their correction relates to infrared behavior, and we are more concerned about the amplitude. Therefore, we ignore this correction for simplicity.,
$A(q)$ $=\frac{1}{2}[f^{\prime}(q)+ic_{s}l(q)],$ (4)
$f(q)$ $=\frac{4\pi}{q}\int_{r_{f}}^{\infty}\mathrm{d}ru(r)\mathrm{sin}(qr),$ (5)
$l(q)$ $=\frac{3\pi}{q}\int_{r_{f}}^{\infty}\mathrm{d}r\delta(r)r\mathrm{sin}(qr).$ (6)
Here $u(r)$ and $\delta(r)$ are the profiles when the sound shells start to collide, $r_{f}$ denotes the cutoff scale since the very high-density perturbation inside and around the PBH will not propagate outward and we choose $r_{f}=20r_{\mathrm{PBH}}$ for a typical illustration. Other choices of $r_{f}$ only slightly change the turning-point position in the power spectrum, where the power law changes roughly from $k^{-9}$ to $k^{-3}$) as shown in the Supplementary Material. As for the sub-critical case, we simply take $r_{f}=0$. The ensemble average of the velocity power spectrum is taken over all sound shells centered at different positions $\vec{x}^{(m)}$ so that
$\sum_{m=1}^{N}\sum_{n=1}^{N}\langle e^{-i\vec{q}\cdot\vec{x}^{(m)}+i\vec{k}\cdot\vec{x}^{(n)}}\rangle=(2\pi)^{3}\delta^{3}(\vec{q}-\vec{k})\frac{N}{V_{c}},$ (7)
where $V_{c}$ is the comoving volume of space, and we can define the comoving mean separation of the sound shells in terms of $R_{*c}^{-3}\equiv N/V_{c}$. Finally, we arrived at
$\langle v_{\vec{q}}^{i}v_{\vec{k}}^{j*}\rangle=\hat{q}^{i}\hat{q}^{j}|A(q)|^{2}(2\pi)^{3}\delta^{3}(\vec{q}-\vec{k})\frac{N}{V_{c}}.$ (8)
For later convenience, we define the dimensionless velocity power spectrum as
$\mathcal{P}_{v}$ $\equiv 2\left(\frac{R_{*c}}{r_{H}}\right)^{3}\frac{q^{3}}{2\pi^{2}R_{*c}^{3}}|A(q)|^{2}=\frac{(qr_{H})^{3}}{\pi^{2}}|A(qr_{H})|^{2},$ (9)
then the GW spectrum follows similarly as Ref. Guo et al. (2021),
$\mathcal{P}_{\mathrm{GW}}(y,kr_{H})=$ $3\Gamma^{2}(H_{s}a_{s}R_{*c})\left(\frac{r_{H}}{R_{*c}}\right)^{7}\frac{(kr_{H})^{3}}{2\pi^{2}}$
×$\times\tilde{P}_{\mathrm{GW}}(kr_{H})\Upsilon(y),$ (10)
where $y$ is a time variable defined as $y=a/a_{s}$, $\Gamma=(\bar{\rho}+\bar{p})/\bar{\rho}\approx 4/3$, the subscript $s$ denotes the time when sound wave starts to source the GWs, roughly the time of sound shell collision, $\Upsilon$ is the usual suppression factor due to Hubble expansion Guo et al. (2021) that approximates to 1 at radiation dominated era and 2/3 at matter-dominated era in the long duration limit, and $\tilde{P}_{\mathrm{GW}}$ is defined as
$\tilde{P}_{\mathrm{GW}}$ $(kr_{H})=\frac{1}{4\pi c_{s}kr_{H}}\left(\frac{1-c_{s}^{2}}{c_{s}^{2}}\right)^{2}$ (11)
×$\times\int_{z_{-}}^{z_{+}}\frac{\mathrm{d}z}{z}\frac{(z-z_{+})^{2}(z-z_{-})^{2}}{z_{+}+z_{-}-z}\bar{P}_{v}(z)\bar{P}_{v}(z_{+}+z_{-}-z)$
[…]
Observational perspectives. As shown in the Supplementary Material, the peak frequency in the GW spectrum is about a quarter Ning et al. (2025) of comoving thickness $d_{c}$ of underdense shell at collisions so that we can define a nearly constant $m\equiv d_{c}f_{\mathrm{peak}}\simeq\mathcal{O}(1)$. On the other hand, the sound shell model has found that the peak frequency is roughly the sound-shell thickness at free propagation, which has been found in our accompanying long paper Ning et al. (2025) that $d_{c}$ will approach a constant proportionality to the comoving Hubble horizon $\mathcal{H}_{f}$ at reentry of initial superhorizon perturbations. Therefore, after characterizing $d_{c}\equiv n\mathcal{H}_{f}^{-1}$ with $n\simeq\mathcal{O}(1)$, we can estimate the present-day peak frequency $f_{0}=(a_{f}/a_{0})f_{\mathrm{peak}}=(a_{f}/a_{0})(m/n)(H_{f}/2\pi)$ in terms of PBH mass $M_{\mathrm{PBH}}=4\pi\gamma M_{\mathrm{Pl}}^{2}/H_{f}$ Carr and Hawking (1974); Hawking (1971); Sasaki et al. (2018) as
$f_{0}\simeq 1.105\times 10^{-2}\,\mathrm{Hz}$ $\left(\frac{106.75}{g_{s}(T_{f})}\right)^{\frac{1}{12}}\left(\frac{g_{s}(T_{0})}{3.94}\right)^{\frac{1}{3}}\left(\frac{\gamma}{0.2}\right)^{\frac{1}{2}}$
×$\times\left(\frac{10^{20}\mathrm{g}}{M_{\mathrm{PBH}}}\right)^{\frac{1}{2}}\left(\frac{4}{n}\right)\left(\frac{m}{4}\right).$ (12)
Therefore, for PBHs of asteroid mass $M_{\mathrm{PBH}}\simeq 10^{20}\,\mathrm{g}$ within the currently open window $(10^{17}-10^{22}\,\mathrm{g})$ with maximally allowable PBH abundance $f_{\mathrm{PBH}}=1$, the corresponding peak frequency $f_{0}\simeq 1.1\times 10^{-2}\,\mathrm{Hz}$ is within the LISA/Taiji/TianQin-DECIGO/BBO frequency range; while for extremely light PBHs of mass $M_{\mathrm{PBH}}<10^{10}\,\mathrm{g}$ that are currently unconstrained, the corresponding peak frequency $f_{0}\simeq 1.1\times 10^{3}\mathrm{Hz}$ is exactly out of the reach of LIGO-Virgo-KAGRA frequency range, that is to say, future GW detectors sensitive to higher frequencies Aggarwal et al. (2021) would, in turn, constrain these extremely light PBHs of mass $M_{\mathrm{PBH}}=10^{-5}-10^{9}\,\mathrm{g}$ that would emit Hawking radiation of particles beyond the current standard model of particle physics. For slightly heavier PBHs of mass $M_{\mathrm{PBH}}=10^{10}-10^{13}\,\mathrm{g}$ constrained by Big Bang nucleosynthesis, the corresponding peak frequency $f_{0}\sim 10-100\,\mathrm{Hz}$ is fully within the LIGO-Virgo-KAGRA frequency range. As for the sub-critical case, one can similarly obtain
$f_{0}\simeq 1.20\times 10^{6}$ $\,\mathrm{Hz}\left(\frac{106.75}{g_{s}(T_{f})}\right)^{\frac{1}{12}}\left(\frac{g_{s}(T_{0})}{3.94}\right)^{\frac{1}{3}}$
×$\times\left(\frac{H_{f}}{0.05\mathrm{Mpc}^{-1}}\right)^{\frac{1}{2}}\left(\frac{4}{n}\right)\left(\frac{m}{4}\right).$ (13)
The amplitude is highly correlated to the mean separation $R_{*c}$ of the sound shell, as there is a factor $(r_{H}/R_{*c})^{7}$ in $\mathcal{P}_{\mathrm{GW}}$. The GW energy spectrum today is expressed as
$h^{2}\Omega_{\mathrm{GW}}(f)=1.66\times 10^{-5}\left(\frac{100}{g_{s}(T_{e})}\right)^{1/3}\mathcal{P}_{\mathrm{GW}}(a_{e}f/a_{0}),$ (14)
with the subscript $e$ for the time of GW production. In the right panel of Fig. 3, we show the present GW energy density spectrum with collision time $t_{c}=50t_{m}$, $t_{c}=100t_{m}$, $t_{c}=200t_{m}$, $t_{c}=400t_{m}$, and $t_{c}=800t_{m}$, respectively. The comoving mean separation of the sound shell is roughly $R_{*c}=134r_{H}$, $R_{*c}=166r_{H}$, $R_{*c}=200r_{H}$, $R_{*c}=280r_{H}$, and $R_{*c}=380r_{H}$. As one can see, a larger $R_{*c}$ or a smaller $\delta_{m}$ is usually correlated to a smaller amplitude of the GW energy spectrum.
Constraints on PBH abundance. One application of this new source of SGWBs is to constrain the abundance of PBHs, especially those that have already evaporated today (see also Ref. Franciolini and Pani (2023)). To investigate the PBH abundance, it is more convenient to use the density fraction of PBHs at the epoch of their formation,
$\beta\equiv\left.\frac{\rho_{\mathrm{PBH}}}{\rho_{\mathrm{FRW}}}\right|_{a=a_{\mathrm{f}}}=\frac{n_{\mathrm{PBH}}M_{\mathrm{PBH},i}}{n_{\mathrm{H}}M_{\mathrm{H}}}=\frac{\gamma}{R_{*}^{3}H_{f}^{3}},$ (15)
where the subscript $f$ denotes the time when the perturbation re-enters the horizon, and $M_{\mathrm{PBH},i}$ is the initial mass of PBHs given by $M_{\mathrm{PBH},i}=\gamma M_{\mathrm{H}}|_{a_{f}}$. In our simulation, the collapsing efficiency $\gamma\simeq 0.355$ and $\mathcal{H}_{f}^{-1}=10r_{H}$. Therefore, we can obtain a simple relation between $\beta$ and $R_{*c}$, $\beta=355(r_{H}/R_{*c})^{3}$, so that $t_{c}=200t_{m}$ corresponds to $\beta=4.4\times 10^{-5}$, $t_{c}=400t_{m}$ corresponds to $\beta=1.6\times 10^{-5}$, and $t_{c}=800t_{m}$ corresponds to $\beta=6.5\times 10^{-6}$. In the right panel of Fig. 3, we have set $M_{\mathrm{PBH}}\simeq 10^{8}\,\mathrm{g}$ in the $\delta_{m}=0.5$ case and $M_{\mathrm{PBH}}\simeq 10^{4}\,\mathrm{g}$ in the $\delta_{m}=0.55$ case. If future observations reveal null detection for the corresponding high-frequency GWs, then the PBH abundance $\beta$ could be constrained to be smaller than $\mathcal{O}(10^{-5})$.
Conclusions and discussions. It is of essential importance to thoroughly study all possible GW sources so as to fully understand the SGWBs in GW detectors. In this Letter, we have proposed a new SGWB sourced by the collisions of sound shells from density perturbation collapse, providing an alternative probe into the evaporated PBH mass and abundance. We note here that a previous work De Luca et al. (2019) has also studied the GW production but during the non-spherical collapses of sub-critical scalar perturbations, which is overwhelmed by those scalar-induced GWs (SIGW) at second-order in perturbations Ananda et al. (2007); Baumann et al. (2007) (see Ref. Domènech (2021) for a recent review and references therein). We stress here that our newly proposed SGWB is complementary to both of the above two GW sources, as the SIGW is mainly produced at horizon crossing, while our SGWB is generated long after horizon crossing when sound waves from gravitational collapses are in full superpositions.
Several aspects of this work could be further explored in future studies: (i) Our evaluation is based on monochromatic density perturbations, assuming that all sound shells are uniformly distributed and formed at roughly the same time. It is natural to consider a spatial or mass distribution, as there might be some cluster effects, which may contribute to a higher amplitude; (ii) As most of our evaluation is numerical, we expect an analytical or semi-analytical profile of sound waves from density perturbation collapse, for this can give us a more general analysis of the final GW energy spectrum; (iii) Our theoretical frame can be applied to other cases where there is a sound wave structure, like PBHs formed from primordial true vacuum bubbles that later become false vacuum bubbles in the radiation era Deng et al. (2018); (iv) As one can see in Fig. 2, the sound wave cannot propagate freely when the separation of the sound shells $R_{*c}$ is too small, which should be considered separately; (v) Not only the sound shell-sound shell collisions but also the sound shell-PBH collisions will generate GWs similar to a recent numerical simulation Yuwen et al. (2024) of bubble wall-PBH collisions.
Acknowledgements.
Acknowledgements. This work is supported by the National Key Research and Development Program of China Grant No. 2021YFC2203004, No. 2021YFA0718304, and No. 2020YFC2201501, the National Natural Science Foundation of China Grants No. 12422502, No. 12105344, No. 12235019, No. 12047503, No. 12073088, No. 11821505, No. 11991052, and No. 11947302, and the Science Research Grants from the China Manned Space Project with No. CMS-CSST-2021-B01 (supported by China Manned Space Program through its Space Application System). HD was supported by Yuri Levin’s Simons Investigator Grant PG012519.
Supplemental Material: Dimensionless gravitational-wave power spectrum from primordial gravitational collapse
Xiang-Xi Zeng1,2 Zhuan Ning,3,1,4 Zi-Yan Yuwen,1,2,5
Shao-Jiang Wang,1,6 Heling Deng,7 and Rong-Gen Cai8
1Institute of Theoretical Physics, Chinese Academy of Sciences (CAS), Beijing 100190, China
2School of Physical Sciences, University of Chinese Academy of Sciences (UCAS), Beijing 100049, China
3School of Fundamental Physics and Mathematical Sciences, Hangzhou Institute for Advanced Study (HIAS), University of Chinese Academy of Sciences (UCAS), Hangzhou 310024, China
4University of Chinese Academy of Sciences (UCAS), Beijing 100049, China
5Department of Physics, Stellenbosch University, Matieland 7602, South Africa
6Asia Pacific Center for Theoretical Physics (APCTP), Pohang 37673, Korea
7Department of Astronomy, Columbia University, New York, NY 10027, USA
8Institute of Fundamental Physics and Quantum Technology, Ningbo University, Ningbo, 315211, China
Appendix A Abstract
This Supplemental Material is prepared for the technical details of the dimensionless gravitational-wave (GW) power spectrum from the primordial gravitational collapse and its comparison to the scalar-induced GWs at second-order in perturbations. The details for implementing long-time numerical simulations of primordial gravitational collapses of density perturbations are prepared in a separate accompanying paper Ning et al. (2025).
Appendix B Appendix A: GW power spectrum from the sound shell model
In this appendix, we closely follow the original sound shell model Hindmarsh (2018); Hindmarsh and Hijazi (2019); Guo et al. (2021). Neglecting the irrelevant scalar and vector perturbations, the metric in the FLRW Universe is defined as:
$\mathrm{d}s^{2}=-\mathrm{d}t^{2}+a(t)^{2}(\delta_{ij}+h_{ij}(t,\vec{x}))\mathrm{d}\vec{x}^{2},$ (16)
where the $h_{ij}$ denotes the transverse traceless tensor perturbation. The GWs are sourced by the transverse traceless part of the energy-momentum tensor $a^{2}\pi^{\mathrm{TT}}$ via
$h_{ij}^{\prime\prime}(t,\vec{k})+2\mathcal{H}h_{ij}^{\prime}(t,\vec{k})+k^{2}h_{ij}(t,\vec{k})=16\pi Ga^{2}\pi_{ij}^{\mathrm{TT}}(t,\vec{k}),$ (17)
where $\mathcal{H}=a^{\prime}/a$, $\prime\equiv\partial/\partial\eta$ and $\eta$ is the conformal time. The GW energy density $\rho_{\mathrm{GW}}$ is defined as
$\rho_{\mathrm{GW}}(t)=\frac{1}{32\pi G}\langle\dot{h}_{ij}(t,\vec{x})\dot{h}_{ij}(t,\vec{x})\rangle.$ (18)
Here, the dot denotes the derivative with respect to the coordinate time. Then we define the power spectrum of $\dot{h}_{ij}$ as
$\langle\dot{h}_{ij}(t,\vec{k})\dot{h}_{ij}(t,\vec{q})\rangle=(2\pi)^{3}\delta^{3}(\vec{k}+\vec{q})P_{\dot{h}}(t,k),$ (19)
and the dimensionless GW energy spectrum follows
$\mathcal{P}_{\mathrm{GW}}(t,k)\equiv\frac{\mathrm{d}\Omega_{\mathrm{GW}}(t)}{\mathrm{d}\ln k}=\frac{1}{\rho_{c}(t)}\frac{\mathrm{d}\rho_{\mathrm{GW}}(t)}{\mathrm{d}\ln k}=\frac{k^{3}}{24\pi^{2}H^{2}}P_{\dot{h}}(t,k)=\frac{k^{3}}{24\pi^{2}H^{2}a^{2}}P_{h^{\prime}}(\eta,k),$ (20)
where the $\rho_{c}$ is the critical energy density at time $t$. To solve the equation (17), one can use the Green function method by introducing a Green function $G(x,x_{0})$ with the following boundary conditions,
$G(x\leq x_{0})=0,~{}~{}~{}~{}~{}\left.\frac{\partial G(x,x_{0})}{\partial x}\right|_{x=x_{0}^{+}}=1,$ (21)
where $x=k\eta$ and $x_{0}$ is the time when the sound wave starts to propagate. Using the Green function, one can obtain
$h_{ij}(t,\vec{k})=16\pi G\int_{x_{0}}^{x}\mathrm{d}x_{1}G(x,x_{1})\frac{a^{2}(\eta_{1})\pi^{\mathrm{TT}}_{ij}(\eta_{1},\vec{k})}{k^{2}},$ (22)
and the 2-point correlation function is expressed as
$\langle h^{\prime}_{ij}(\eta,\vec{k})h^{\prime}_{ij}(\eta,\vec{q})\rangle=(16\pi G)^{2}\int_{x_{0}}^{x}\mathrm{d}x_{1}\int_{x_{0}}^{x}\mathrm{d}x_{2}\frac{\partial G(x,x_{1})}{\partial x}\frac{\partial G(x,x_{2})}{\partial x}\frac{a^{2}(\eta_{1})a^{2}(\eta_{2})}{q^{2}}\langle\pi^{\mathrm{TT}}_{ij}(\eta_{1},\vec{k})\pi^{\mathrm{TT}}_{ij}(\eta_{2},\vec{q})\rangle.$ (23)
Therefore, one should evaluate the unequal time correlator (UETC) of $\pi_{ij}^{\mathrm{TT}}$. Due to the spatial homogeneity of the Universe, one can parameterize it as
$\langle\pi^{\mathrm{TT}}_{ij}(\eta_{1},\vec{k})\pi^{\mathrm{TT}}_{ij}(\eta_{2},\vec{q})\rangle=\Pi^{2}(k,\eta_{1},\eta_{2})(2\pi)^{2}\delta^{3}(\vec{k}+\vec{q}).$ (24)
The only contributing part of $\pi_{ij}^{\mathrm{TT}}$ is
$\pi_{ij}^{\mathrm{TT}}=(p+\rho)\gamma^{2}v^{i}v^{j},$ (25)
where $p$ is the pressure, $\rho$ is the energy density, $v^{i}$ is the velocity defined by $v^{i}=\mathrm{d}x^{i}/\mathrm{d}\eta$, and $\gamma$ is the Lorentz factor that can be taken as 1 in the non-relativistic case. As for the velocity field, we use $\tilde{v}_{\vec{k}}^{i}$ to denote the Fourier transformation of $v^{i}(\vec{x})$. Neglecting the rotational component Hindmarsh and Hijazi (2019), the 2-point correlator can be defined as
$\langle\tilde{v}^{i}_{\vec{k}}(\eta_{1})\tilde{v}^{j*}_{\vec{q}}(\eta_{2})\rangle=(2\pi)^{3}\delta^{3}(\vec{q}-\vec{k})\hat{q}^{i}\hat{k}^{j}B(q,\eta_{1},\eta_{2}),$ (26)
then one can directly calculate the UETC of $\pi_{ij}^{\mathrm{TT}}$ and find
$\Pi^{2}(k,\eta_{1},\eta_{2})=\bar{\omega}^{2}\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}B(q,\eta_{1},\eta_{2})B(\tilde{q},\eta_{1},\eta_{2})\frac{q^{2}}{\tilde{q}^{2}}(1-\mu^{2})^{2},$ (27)
where $\bar{\omega}=\bar{\rho}+\bar{p}$ denotes the background’s homogeneous value, $\tilde{q}=|\vec{q}-\vec{k}|$ and $\mu=\hat{q}\cdot\hat{k}$. The velocity field can be Fourier decomposed as
$v^{i}(\eta,\vec{x})=\frac{1}{2}\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}[\tilde{v}^{i}_{\vec{q}}(\eta)e^{i\vec{q}\cdot\vec{x}}+\tilde{v}^{i*}_{\vec{q}}(\eta)e^{-i\vec{q}\cdot\vec{x}}],$ (28)
where $\vec{x}$ and $\vec{q}$ are both comoving quantities. If the velocity field propagates freely as sound waves after some initial time $\eta_{i}$, it will also admit the plane wave solutions
$v^{i}(\eta>\eta_{i},\vec{x})=\int\frac{\mathrm{d}^{3}q}{(2\pi)^{3}}[v^{i}_{\vec{q}}e^{-i\omega\eta+i\vec{q}\cdot\vec{x}}+v^{i*}_{\vec{q}}e^{i\omega\eta-i\vec{q}\cdot\vec{x}}],$ (29)
where $\omega=c_{s}q$, $c_{s}=1/\sqrt{3}$ is the sound speed, and for the longitudinal plane wave, we have $v_{\vec{q}}^{i}=v_{\vec{q}}\hat{q}^{i}$. Note that $v^{i}_{\vec{q}}$ is the amplitude of the plane wave solution, which is different from $\tilde{v}_{\vec{q}}^{i}$. Then one can express the $B(q,\eta_{1},\eta_{2})$ in terms of the 2-point function of $v_{\vec{q}}$ in the stationary case Hindmarsh and Hijazi (2019); Roper Pol et al. (2024),
$B(q,\eta_{1},\eta_{2})=2P_{v}(q)\cos[\omega(\eta_{1}-\eta_{2})],$ (30)
where $\langle v_{\vec{q}}v_{\vec{k}}^{*}\rangle=P_{v}(q)(2\pi)^{3}\delta^{3}(\vec{k}-\vec{q})$. After defining the energy fluctuation
$\lambda(x)=\frac{\rho(x)-\bar{\rho}}{\bar{\omega}},$ (31)
the equations for sound waves follow
$\tilde{\lambda}_{\vec{q}}^{\prime}+iq^{j}\tilde{v}^{j}_{\vec{q}}=0,$ (32)
$\tilde{v}_{j}^{\prime}+c_{s}iq^{j}\tilde{\lambda}_{\vec{q}}=0.$ (33)
Using these equations, one can find at some initial time $\eta_{i}$, the plane wave amplitude is Hindmarsh and Hijazi (2019)
$v_{\vec{q}}=\frac{1}{2}\left(\hat{q}^{i}\tilde{v}^{i}_{\vec{q}}(\eta_{i})-c_{s}\tilde{\lambda}_{\vec{q}}(\eta_{i})\right)e^{i\omega\eta_{i}}.$ (34)
Therefore, one should assign the initial condition for the $\tilde{v}^{i}_{\vec{q}}(\eta_{i})$ and $\tilde{\lambda}_{\vec{q}}(\eta_{i})$. In coordinate space, the velocity profile for the $n$-th sound shell created by curvature perturbation can be written as
$v_{i}^{(n)}(\eta,\vec{x})=\frac{R_{i}^{(n)}}{R^{(n)}}u(\eta,R^{(n)}),$ (35)
where $\vec{R}^{(n)}=\vec{x}-\vec{x}^{(n)}$, $\vec{x}^{(n)}$ is the center of the $n$-th sound shell and $u(\eta,r)$ is radial profile of the velocity field. Assuming $\bar{\omega}=\Gamma\bar{\rho}$, one can write the energy fluctuation as
$\lambda^{(n)}(\eta,\vec{x})=\frac{1}{\Gamma}\delta^{(n)}(\eta,\vec{x})=\frac{1}{\Gamma}\delta(\eta,R^{(n)}),$ (36)
where $\delta(\eta,R^{(n)})$ denotes the profile of density contrast. After Fourier transformation, one can obtain
$\tilde{v}^{i(n)}_{\vec{q}}(\eta)=e^{-i\vec{q}\cdot\vec{x}^{(n)}}i\hat{q}^{i}f^{\prime}(\eta,q),$ (37)
$\tilde{\lambda}^{(n)}_{\vec{q}}(\eta)=e^{-i\vec{q}\cdot\vec{x}^{(n)}}l(\eta,q),$ (38)
and functions $f(\eta,q),l(\eta,q)$ are given by
$f(\eta,q)=\frac{4\pi}{q}\int_{0}^{\infty}\mathrm{d}ru(\eta,r)\sin(qr),$ (39)
$l(\eta,q)=\frac{4\pi}{\Gamma q}\int_{0}^{\infty}\mathrm{d}r\delta(\eta,r)r\sin(qr).$ (40)
It is convenient to define a function $A(\eta,r)\equiv(f^{\prime}(\eta,r)+ic_{s}l(\eta,r))/2$, then we have
$v_{\vec{q}}^{(n)}=ie^{i\omega\eta_{i}-i\vec{q}\cdot\vec{x}^{(n)}}A(\eta_{i},q).$ (41)
Now we can calculate the spectral density of the plane wave amplitudes, assuming all the sound shells are propagating freely at roughly the same time $\eta_{i}$,
$\langle v_{\vec{q}}^{i}v_{\vec{k}}^{j*}\rangle=\sum_{m=1}^{N}\sum_{n=1}^{N}\hat{q}^{i}\hat{k}^{j}A(q)A^{*}(k)\langle e^{-i\vec{q}\cdot\vec{x}^{(m)}+i\vec{k}\cdot\vec{x}^{(n)}}\rangle e^{i(\omega_{q}-\omega_{k})\eta_{i}},$ (42)
with $N$ denoting the number of the sound shell, then the ensemble average gives rise to
$\sum_{m=1}^{N}\sum_{n=1}^{N}\langle e^{-i\vec{q}\cdot\vec{x}^{(m)}+i\vec{k}\cdot\vec{x}^{(n)}}\rangle=(2\pi)^{3}\delta^{3}(\vec{q}-\vec{k})\frac{N}{V_{c}},$ (43)
where $V_{c}$ is the comoving volume of space. Next we can define the comoving mean separation of the sound shell in terms of $R_{*c}^{-3}\equiv N/V_{c}$. Finally, we arrived at
$\langle v_{\vec{q}}^{i}v_{\vec{k}}^{j*}\rangle=\hat{q}^{i}\hat{q}^{j}|A(q)|^{2}(2\pi)^{3}\delta^{3}(\vec{q}-\vec{k})\frac{N}{V_{c}}.$ (44)
The dimensionless velocity power spectrum is defined as [^3] In our main text, we have defined $\mathcal{P}_{v}\equiv 2\left(\frac{R_{*c}}{r_{H}}\right)^{3}\frac{q^{3}}{2\pi^{2}R_{*c}^{3}}|A(q)|^{2}$ since this is more convenient to see the behavior of $\frac{(kr_{H})^{3}}{2\pi^{2}}\tilde{P}_{\mathrm{GW}}(kr_{H})$ without the influence of $R_{*c}$.
$\mathcal{P}_{v}(q)\equiv\frac{q^{3}}{2\pi^{2}}[2P_{v}(q)]=\frac{q^{3}}{\pi^{2}R_{*c}^{3}}|A(q)|^{2},$ (45)
where the factor 2 comes from (30). After solving the Green function Guo et al. (2021), for example, the Green function in RD dominated era  [^4] For the general solution, one can refer to Ref. Guo et al. (2021). is
$G(x,x_{0})=\frac{x_{0}\sin(x-x_{0})}{x},$ (46)
then one can derive the final dimensionless GW energy spectrum
$\mathcal{P}_{\mathrm{GW}}(y,kr_{H})=3\Gamma^{2}(H_{s}a_{s}r_{H})\frac{(kr_{H})^{3}}{2\pi^{2}}\tilde{P}_{\mathrm{GW}}(kr_{H})\Upsilon(y),$ (47)
where $y$ is a time variable defined as $y=a/a_{s}$, $r_{H}$ is some characteristic length scale, the subscript $s$ denotes the time when the sound wave starts to source the GWs, roughly the time of sound shell collision, $\Upsilon$ is the usual suppression factor due to Hubble expansion Guo et al. (2021) that approximates to 1 at radiation dominated era and 2/3 at matter-dominated era in the long duration limit, and $\tilde{P}_{\mathrm{GW}}$ is defined as
$\tilde{P}_{\mathrm{GW}}(kr_{H})=\frac{1}{4\pi c_{s}kr_{H}}\left(\frac{1-c_{s}^{2}}{c_{s}^{2}}\right)^{2}\int_{z_{-}}^{z_{+}}\frac{\mathrm{d}z}{z}\frac{(z-z_{+})^{2}(z-z_{-})^{2}}{z_{+}+z_{-}-z}\bar{P}_{v}(z)\bar{P}_{v}(z_{+}+z_{-}-z).$ (48)
with $z=qr_{H}$, $z_{\pm}=\frac{kr_{H}}{2c_{s}}(1\pm c_{s})$, and $\bar{P}(z)=\frac{\pi^{2}\mathcal{P}_{v}(z)}{z^{3}}$.
Appendix C Appendix B: The dimensionless GW power spectrum with different cutoff scale $r_{f}$
As one can see in Fig. 4, the peak amplitude of the dimensionless power spectrum is not sensitive to the choice of cutoff scale $r_{f}$, which only slightly affects the broken power law position where the power law changes from roughly $k^{-9}$ to $k^{-3}$.
Figure 4: The dimensionless GW power spectrum with different cutoff scale $r_{f}$.
Appendix D Appendix C: The peak behavior of the GW power spectrum
Similar to Ref. Hindmarsh (2018), we can analyze the peak position and power law of the final GW power spectrum. Starting from the super-critical case. We use a test function,
$u(r)=\frac{a}{l}(r-r_{0})\Theta(r-r_{0}+l)\Theta(r_{0}-r),$ (49)
to characterize the key features of the sound shell profile, for example, the position $r_{0}$ of the wavefront, the thickness $l$ (roughly at the same order of the sound-shell thickness $d_{c}$ when propagating freely until collision), and depth $a$ of the underdense region, where $\Theta$ is the step function and the sub-dominated overdense region is neglected for simplicity. We will use the approximated relation  [^5] In our numerical calculation, we will directly use the initial profile of $\delta(r)$ from numerical simulations, and we will only use this approximated relation in the analytical analysis in the Supplementary Material. $\delta(r)=4c_{s}u(r)$ found in Ref. Deng (2021), which is consistent with our numerical results. Here are two examples shown in Fig. 5.
Figure 5: Left column: The test function with $a=0.03,l=40,r_{0}=144$. Right column: The test function with $a=0.02,l=40,r_{0}=190$. The top panels show the velocity profile, while the bottom panels show the dimensionless GW power spectrum.
For the velocity power spectrum $\mathcal{P}_{v}$ of a form $(qr_{H})^{n}$ (9) over a range of wave numbers, the GW power spectrum $\mathcal{P}_{\mathrm{GW}}(kr_{H})$ becomes $(kr_{H})^{2n-1}$. After some algebra, one can obtain
$f^{\prime}(q)=\frac{4\pi a}{lq^{4}}\bigg{(}qr_{0}\cos(qr_{0})+q(3l-r_{0})\cos(q(r_{0}-l))-3\sin(qr_{0})+(3+lq^{2}(r_{0}-l))\sin(q(r_{0}-l))\bigg{)},$ (50)
$l(q)=\frac{4\sqrt{3}\pi a}{lq^{4}}\bigg{(}2\cos(qr_{0})+(lq^{2}(l-r_{0})-2)\cos(q(r_{0}-l))+qr_{0}\sin(qr_{0})+q(2l-r_{0})\sin(q(r_{0}-l))\bigg{)},$ (51)
then we can calculate $\mathcal{P}_{v}$ at different limits,
$\mathcal{P}_{v}(qr_{H})=\left\{\begin{aligned} &\frac{a^{2}(qr_{H})^{3}}{4}\left(\frac{(qr_{H})^{2}}{225}(15l^{3}r_{0}-20l^{2}r_{0}^{2}+10lr_{0}^{3}-4l^{4})^{2}+\frac{1}{36}(16l^{2}r_{0}-12lr^{2}_{0}-6l^{3})^{2}\right),~{}~{}~{}~{}~{}ql,qr_{0}\ll 1;\\ &a^{2}l^{2}r_{0}^{2}qr_{H},~{}~{}~{}~{}~{}ql\ll 1\ll qr_{0};\\ &4a^{2}(r_{0}-l)^{2}(qr_{H})^{-1},~{}~{}~{}~{}1\ll ql,qr_{0}.\end{aligned}\right.$
finally, we recover the power law behavior of $\mathcal{P}_{\mathrm{GW}}(kr_{H})$ as
$\mathcal{P}_{\mathrm{GW}}(kr_{H})\sim\left\{\begin{aligned} &(kr_{H})^{5},~{}~{}~{}~{}~{}kl,kr_{0}\ll 1,\\ &(kr_{H})^{1},~{}~{}~{}~{}~{}kl\ll 1\ll kr_{0},\\ &(kr_{H})^{-3},~{}~{}~{}~{}1\ll kl,kr_{0}.\end{aligned}\right.$
It is worth noting that, in some special cases, the $(qr_{H})^{5}$ term may dominate $\mathcal{P}_{v}$ when $ql,qr_{0}\ll 1$, which will contribute a $(kr_{H})^{9}$ term in $\mathcal{P}_{\mathrm{GW}}(kr_{H})$. Note also that, in the domain $ql\ll 1\ll qr_{0}$, we cannot ignore the term like $2\pi a(ql)^{2}\sin(qr_{0})/(q^{4}l)$ in $f^{\prime}(q)$ when $r_{0}$ is not much bigger than $l$, since this term will slightly change the power law within the range of $(qr_{H})^{-1}\sim(qr_{H})^{1}$, which will lead to a change in the power law of $\mathcal{P}_{\mathrm{GW}}(kr_{H})$ within the range of $(kr_{H})^{-3}\sim(kr_{H})^{1}$. Equipped with the above results, one can easily find that the higher peak frequency is related to $l$, and the lower peak frequency is related to $r_{0}$. The analysis of near-critical cases and sub-critical cases is the same but much more complicated. However, it is not difficult to find that the higher peak frequency is related to the thickness of the overdense region, for this is the smallest scale in $u(r)$. Here, the thicknesses of the overdense and underdense regions are roughly the same; therefore, we can use the thickness of the underdense region to determine the peak frequency to be consistent with the super-critical cases.
Appendix E Appendix D: Comparison with Scalar Induced Gravitational Waves (SIGW)
It is well known that enhanced curvature perturbations can also induce GWs at second-order in perturbations Ananda et al. (2007); Baumann et al. (2007). Therefore, we should compare the amplitude of this scalar-induced GW (SIGW) with our GW generated by sound waves (SWGW). Before diving into the comparison, we should stress that our mechanism applies not only to the case with curvature perturbations but also to other cases that generate sound waves, such as that from Ref. Deng et al. (2018). As a rough estimation, we will use a Gaussian probability distribution function (PDF) for the density perturbations,
$\mathscr{P}_{\mathrm{G}}[\delta]=\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{\delta^{2}}{2\sigma^{2}}},$ (52)
where $\sigma^{2}$ is the variance of the distribution calculated by the power spectrum of the density perturbation as
$\sigma^{2}=\langle\delta^{2}\rangle=\int_{0}^{\infty}\mathrm{d}\ln q\mathcal{W}^{2}(q,R)\mathcal{P}_{\delta}(q).$ (53)
Here, $\mathcal{W}$ denotes the Fourier transform of a real space window function and $R$ is the characteristic scale $R\approx\mathcal{H}^{-1}$. In the following calculations, we will choose a top hat window function for a simple illustration,
$\mathcal{W}(q,R)=\frac{3\sin(qR)-3qR\cos(qR)}{(qR)^{3}}.$ (54)
The relation between the power spectrum of density perturbation and curvature perturbation at linear order is given by Josan et al. (2009)
$\mathcal{P}_{\delta}(k)\simeq\frac{16}{81}\left(\frac{k}{aH}\right)^{4}\mathcal{P}_{\mathcal{R}}(k),$ (55)
which, after combined with (53), can be used to relate $\mathcal{P}_{\mathcal{R}}(k)$ with $\sigma^{2}$. Then, using the Press-Schechter model Press and Schechter (1974), one can estimate $\beta$ in terms of $\sigma$ as
$\beta\equiv\int_{\Delta_{c}}^{\infty}\mathscr{P}[\delta]\mathrm{d}\delta.$ (56)
In the following, we will ignore the actual relation between the density contrast $\Delta_{c}$ and $\delta_{c}$, and we will simply choose $\Delta_{c}=\delta_{m}$ for our cases. For a delta-peak curvature perturbation power spectrum $\mathcal{P}_{\mathcal{R}}=A\delta(\ln(k/k_{*}))$, we find the amplitude $A\simeq 10^{-2}$ for all cases, thus, we will choose $A=10^{-2}$ for our rough estimation. Using a more realistic Gaussian power spectrum,
$\mathcal{P}_{\mathcal{R}}(k)=\frac{A(k/k_{*})^{3}}{\sqrt{2\pi}\epsilon}e^{-\frac{((k/k_{*})-1)^{2}}{2\epsilon^{2}}},$ (57)
and choosing $A=10^{-2},\epsilon=10^{-1},k_{*}=\mathcal{H}_{f}=(10r_{H})^{-1}$ for a fair comparison, we can calculate the corresponding SIGW and SWGWs as shown in Fig. 6. As one can see, besides the slightly different peak frequencies, although the amplitude of SIGW is much larger than the corresponding SWGWs at low frequencies, the high-frequency part of SWGWs might be detectable, as there is a high-frequency cutoff in SIGW. We further note here that, although the SIGW seems to be quite similar to our SWGW in nature, as they are all generated from the scalar perturbations in the early Universe, their production eras are different: the SIGW is mainly produced at the time of horizon re-entry, while our SWGW is mainly produced long after horizon re-entry when the sound waves from gravitational collapses are in full superpositions.
Figure 6: The GW energy density spectra of scalar-induced gravitational waves (pink) and our gravitational waves from sound waves of primordial gravitational collapses.